# Structured Agent Distillation for Large Language Model

### v2

@@ A.2 Reasoning–Action Segmentation
A.3 Loss Flow in Structured Agent Distillation
Figure 5: An overview of the loss flow in our Structured Agent Distillation framework. The student model learns from the teacher’s trajectory by decomposing the learning signals into reasoning and action objectives. These loss terms are aggregated to guide effective student optimization.
@@ A.3 Loss Flow in Structured Agent Distillation
A.4 Curriculum Sampling Process
To guide the student model’s learning in a progressive manner, we apply a curriculum learning strategy. The curriculum scheduler $\mathcal{C}$ starts with easier reasoning-action trajectories and gradually introduces harder samples as training progresses. This improves training stability and enhances generalization, especially in long-horizon tasks.
Figure 6: Illustration of the curriculum sampling process. The scheduler gradually increases task difficulty during training, guiding the student model from simple to complex structured agent trajectories under Structured Agent Distillation.
@@ Appendix B Task and Environment Setup
ALFWorld. ALFWorld [26] is an embodied instruction-following environment where agents must complete household tasks through navigation and object manipulation. Each episode provides a natural language goal (e.g., "Put the soap on the sink"), and the agent interacts with a simulated environment to achieve it. We collect teacher trajectories from a ReAct-style agent that interleaves reasoning and actions, and segment them into reasoning spans and action spans.
WebShop. WebShop [27] is a text-based e-commerce environment where agents must fulfill shopping goals through search, click, and buy tool invocations. Teacher agents generate structured trajectories consisting of reasoning steps (e.g., product comparisons) and concrete tool actions. Students are trained on these structured trajectories under curriculum sampling.
HotPotQA-ReAct. We adapt HotPotQA [1] into a ReAct-style question answering setup, where teacher agents interleave multi-hop reasoning steps before final answers. Trajectories are segmented into reasoning and action spans, ensuring a clear separation between intermediate thoughts and conclusive outputs. Students are trained to imitate these structured trajectories under segment-aware supervision.
@@
All models are trained on A6000 GPU, and evaluation is conducted using deterministic greedy decoding unless otherwise noted.
@@ Appendix C Evaluation Metrics
  • Task Success Rate (TSR) [28]: Measures the success ratio of tasks completed by the agent:
@@
  • Latency Results (Average Reasoning Steps) [76]. As a proxy for agent latency, we measure the average number of generation steps required to complete the reasoning and action phases. Given $N$ trajectories:
@@ D.1 Trajectory Segmentation
Table 6: Segmentation rules for splitting teacher trajectories into reasoning and action spans.
@@
D.2 Segment-Aware Token Mask Construction
@@
D.3 Example and Loss Application
@@ E.1 Prompt Templates
E.2 Teacher Trajectory Collection
@@ Appendix F Validation of Segment-Aware Mask Construction
F.1 Mask Construction Rules
@@
F.2 Segment-Aware Mask Visualizations
@@
Figure 7: Segment-Aware Token Mask Overlay across Tasks. We visualize the token-level masks used in Structured Agent Distillation across three environments: WebShop, HotPotQA, and ALFWorld. Reasoning spans are shown in blue, and action spans in red. These overlays validate the alignment between the teacher’s generated structure and our mask segmentation, supporting structure-aware loss formulation.
@@
F.3 Span Length Statistics.
@@
Figure 8: Token Span Length Statistics Across Tasks. Left: Token length distribution for reasoning spans. Right: Token length distribution for action spans. Reasoning spans tend to be longer and more variable, while action spans are concise and more consistent.
@@ Appendix G Additional Experimental Results and Analysis
Table 7: Task success rate (%) comparing OPT-13B and LLaMA-13B teachers, token-level baselines [13], and our Structured Agent Distillation students across model sizes.
@@
Table 8: Average reasoning length (tokens) comparing OPT-13B and LLaMA-13B teachers, token-level baselines [13], and Structured Agent Distillation students.
@@
Table 9: Chain-of-Thought (CoT) match rate (%) comparing OPT-13B and LLaMA-13B teachers, token-level baselines [13], and Structured Agent Distillation students.
@@
Table 10: Latency results (average steps per episode) comparing OPT-13B and LLaMA-13B teachers, token-level baselines [13], and Structured Agent Distillation students. Lower is better.
@@
Figure 9: Scaling analysis on OPT model family. We compare task success rate, reasoning efficiency (avg reasoning length), CoT match rate, and latency across three OPT-based student model sizes (1.3B, 2.7B, 6.7B), using Structured Agent Distillation vs. token-level MiniLLM [13] baselines. The OPT-13B teacher serves as a reference upper bound. Structured Agent Distillation consistently yields stronger performance across all metrics.

### v3

@@ A.2. Reasoning–Action Segmentation
A.3. Loss Flow in Structured Agent Distillation
Figure 6. An overview of the loss flow in our Structured Agent Distillation framework. The student model learns from the teacher’s trajectory by decomposing the learning signals into reasoning and action objectives. These loss terms are aggregated to guide effective student optimization.
[…]
An overview of the loss flow in our Structured Agent Distillation framework. The student model learns from the teacher’s trajectory by decomposing the learning signals into reasoning and action objectives. These loss terms are aggregated to guide effective student optimization.
@@ A.3. Loss Flow in Structured Agent Distillation
A.4. Curriculum Sampling Process
In Figure 7, to guide the student model’s learning in a progressive manner, we apply a curriculum learning strategy. The curriculum scheduler $\mathcal{C}$ starts with easier reasoning-action trajectories and gradually introduces harder samples as training progresses. This improves training stability and enhances generalization, especially in long-horizon tasks.
Figure 7. Curriculum Sampling Process. The scheduler gradually increases task difficulty during training, guiding the student model from simple to complex structured agent trajectories under Structured Agent Distillation.
[…]
Illustration of the curriculum sampling process. The scheduler gradually increases task difficulty during training, guiding the student model from simple to complex structured agent trajectories under Structured Agent Distillation.
@@ Appendix B Task and Environment Setup
ALFWorld. ALFWorld ALFWorld20 is an embodied instruction-following environment where agents must complete household tasks through navigation and object manipulation. Each episode provides a natural language goal (e.g., ”Put the soap on the sink”), and the agent interacts with a simulated environment to achieve it. We collect teacher trajectories from a ReAct-style agent that interleaves reasoning and actions, and segment them into reasoning spans and action spans.
WebShop. WebShop yao2022webshop is a text-based e-commerce environment where agents must fulfill shopping goals through search, click, and buy tool invocations. Teacher agents generate structured trajectories consisting of reasoning steps (e.g., product comparisons) and concrete tool actions. Students are trained on these structured trajectories under curriculum sampling.
HotPotQA-ReAct. We adapt HotPotQA yao2023react into a ReAct-style question answering setup, where teacher agents interleave multi-hop reasoning steps before final answers. Trajectories are segmented into reasoning and action spans, ensuring a clear separation between intermediate thoughts and conclusive outputs. Students are trained to imitate these structured trajectories under segment-aware supervision.
Models. We apply Structured Agent Distillation to compress ReAct-style GPT-2-1.5B teacher trajectories into student models of 120M, 340M, and 760M (GPT-2); OPT-13B into OPT students (1.3B, 2.7B, 6.7B); and LLaMA-13B into LLaMA-7B, Orca2-13B into Orca2-7B.
Baselines. We compare against token-level baselines, primarily using MiniLLM gu2024minillm—a state-of-the-art student model distilled via token-level imitation, which minimizes the KL divergence between student and teacher outputs at every decoding step.KD distilbert; lightpaff fine-tunes the student model on $\mathcal{D}$ using the teacher distribution as the supervision at each token step, also known as word-level KD. SeqKD (skd; vicuna; alpaca; ITGPT4; lima) fine-tunes the student model on the data generated by the teacher model.
@@
To support generalization, we used adapter-based tuning, optional KL regularization, and span-level supervision, which helps reduce overfitting. The details are in Appendix J.
The teacher is queried online to avoid distribution drift: offline trajectories become stale as the student deviates from the teacher’s rollout policy, reducing success rates by 2–3%. In practice, cached teacher logits make online distillation efficient with negligible overhead.
Inference Details We followed standard evaluation practice by running each method on 3 independent random seeds per task, using fixed train/test splits. Results are reported as the mean across runs; standard deviation was negligible and omitted for clarity. All models are trained on A6000 GPU, and evaluation is conducted using deterministic greedy decoding unless otherwise noted.
@@ Appendix C Evaluation Metrics
  • Task Success Rate (TSR) liu2023agentbench: Measures the success ratio of tasks completed by the agent:
@@
    This measures the proportion of teacher reasoning tokens reproduced by the student, averaged over $N$ samples. Alignment is performed at the token level, ensuring robustness to minor length variation but not to paraphrasing.
    Limitations. CoT Match captures lexical overlap rather than semantic equivalence; hence, logically consistent but rephrased reasoning may be undervalued. We discuss this limitation in the revision and note that future extensions (e.g., BERTScore-based semantic matching) can complement the metric.
@@
  • Latency Results (Average Reasoning Steps) jin2024impact. As a proxy for agent latency, we measure the average number of generation steps required to complete the reasoning and action phases. Given $N$ trajectories:
@@ D.1. Trajectory Segmentation
Table 7. Segmentation rules for splitting teacher trajectories into reasoning and action spans.
@@
D.2. Segment-Aware Token Mask Construction
@@
D.3. Example and Loss Application
@@ E.1. Prompt Templates
E.2. Teacher Trajectory Collection
@@ Appendix F Validation of Segment-Aware Mask Construction
F.1. Mask Construction Rules
@@
F.2. Segment-Aware Mask Visualizations
@@
Figure 8. Segment-Aware Token Mask Overlay across Tasks. We visualize the token-level masks used in Structured Agent Distillation across three environments: WebShop, HotPotQA, and ALFWorld. Reasoning spans are shown in blue, and action spans in red. These overlays validate the alignment between the teacher’s generated structure and our mask segmentation, supporting structure-aware loss formulation.
[…]
Segment-Aware Token Mask Overlay across Tasks.
@@
F.3. Span Length Statistics.
@@
The Y-axis (Frequency) represents the normalized histogram count of reasoning or action span lengths across all trajectories within each dataset. Specifically, for each task, we compute the distribution
[…]
so the sum of frequencies equals 1.
The observed token length ranges (reasoning ¡ 20, action ¡ 8) reflect the structure of ReAct-style datasets rather than a limitation of SAD: in ALFWorld and WebShop, most reasoning traces are concise one- or two-sentence thoughts, and actions are short textual commands (e.g., pickup[obj], search[keyword]). Longer reasoning segments occur in HotPotQA-ReAct but are still under 20 tokens after tokenization because the prompts are already decomposed into multi-step sub-questions. Hence, the narrow token count range accurately characterizes the empirical distribution of reasoning and action spans in the evaluated benchmarks.
Figure 9. Token Span Length Statistics Across Tasks. Left: Token length distribution for reasoning spans. Right: Token length distribution for action spans. Reasoning spans tend to be longer and more variable, while action spans are concise and more consistent.
[…]
Token Span Length Statistics Across Tasks.
@@ Appendix G Additional Experimental Results and Analysis
Table 8. Task success rate (%) comparing OPT-13B and LLaMA-13B teachers, token-level baselines gu2024minillm, KD distilbert; lightpaff, SeqKD skd; vicuna; alpaca; ITGPT4; lima methods, and our Structured Agent Distillation students across model sizes.
@@
Table 9. Average reasoning length (tokens) comparing OPT-13B and LLaMA-13B teachers, token-level baselines gu2024minillm, KD distilbert; lightpaff, SeqKD skd; vicuna; alpaca; ITGPT4; lima, and Structured Agent Distillation students.
@@
Table 10. Chain-of-Thought (CoT) match rate (%) comparing OPT-13B and LLaMA-13B teachers, token-level baselines gu2024minillm, KD distilbert; lightpaff, SeqKD skd; vicuna; alpaca; ITGPT4; lima, and Structured Agent Distillation students.
@@
Table 11. Latency results (average steps per episode) comparing OPT-13B and LLaMA-13B teachers, token-level baselines gu2024minillm, KD distilbert; lightpaff, SeqKD skd; vicuna; alpaca; ITGPT4; lima, and Structured Agent Distillation students. Lower is better.
@@
Figure 10. Scaling analysis on OPT model family. We compare task success rate, reasoning efficiency (avg reasoning length), CoT match rate, and latency across three OPT-based student model sizes (1.3B, 2.7B, 6.7B), using Structured Agent Distillation vs. token-level MiniLLM gu2024minillm baselines. The OPT-13B teacher serves as a reference upper bound. Structured Agent Distillation consistently yields stronger performance across all metrics.
[…]
Scaling analysis on OPT model family.